given vector A = (2i+3j+k), vector B(3i+2j+4k) find (A+B)*(A-B)

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
A + B = 5(i + j + k)
A – B = – i + j – 3k
 
cross product
hence (A – B) * (A +B) = 5( k + 3j + 3i + k – j – i) = 10( i + j + k)